Grade 12Integral Calculus

I=xdx/(√7x-10-x2)3

B.V.Suhas Sudheendhra
14 Years agoGrade 12

1 Answer

Chetan Mandayam Nayakar
ApprovedApproved Tutor Answer14 Years ago

7x-10-x2=2.25-(x-3.5)2

let x-3.5=t

I=∫(t+3.5)dt/(√2.25-t2)3=∫tdt/(√2.25-t2)3+∫(3.5)dt/(√2.25-t2)3